Show that (Cp-Cv) d²T/dPdV + (dCp/dP)v (dT/dV)p - (dCv/dV)p (dT/dP)v =1 using maxwell's equation
1
Expert's answer
2020-11-09T06:49:17-0500

I am assuming that gas is ideal and number of moles is 1.

then, by ideal gas equation, I can write it as "PV = RT"


Now,

"(\frac{\partial }{\partial V})_P [(\frac{\partial }{\partial P})_V(PV)] =(\frac{\partial }{\partial V})_P [(\frac{\partial }{\partial P})_V(TR)]"


Since, "C_P - C_V = R"

Then equation will be,

"(\frac{\partial }{\partial V})_P [(\frac{\partial }{\partial P})_V(PV)] =(\frac{\partial }{\partial V})_P [(\frac{\partial }{\partial P})_V[T(C_P-C_V) ]]"


Differentiating both sides with respect to P keeping V as constant,

"(\frac{\partial }{\partial V})_P [(\frac{\partial P}{\partial P})_VV] =(\frac{\partial }{\partial V})_P [(\frac{\partial T }{\partial P})_V(C_P-C_V)+T(\frac{\partial C_P }{\partial P})_V]"

"(\frac{\partial }{\partial V})_P [V] =(\frac{\partial }{\partial V})_P [(\frac{\partial T }{\partial P})_V(C_P-C_V)+T(\frac{\partial C_P }{\partial P})_V]"


Again differentiating with respect to V keeping P as constant,

"(\frac{\partial V}{\partial V})_P =(\frac{\partial^2 T}{\partial V \partial P})(C_P-C_V)+(\frac{\partial C_P }{\partial P})_V (\frac{\partial T}{\partial V})_P -(\frac{\partial C_V }{\partial V})_P(\frac{\partial T}{\partial P})_V"


It can be written as,

"1 =(\frac{\partial^2 T}{\partial V \partial P})(C_P-C_V)+(\frac{\partial C_P }{\partial P})_V (\frac{\partial T}{\partial V})_P -(\frac{\partial C_V }{\partial V})_P(\frac{\partial T}{\partial P})_V"

which is the required result.